# Eaman Jahani

### v5

@@ A.2.1 Generative Models
    DALL-E 2, which is referred to at points in the main text as the inferior model.
@@
    DALL-E 3 (Verbatim), which is referred to at points in the main text as the superior model.
@@
    DALL-E 3 (Revised), which is referred to at points in the main text as DALL-E 3 with Revision
Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API. OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E image-generating model. During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system. To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3. This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^8] See here and here for online documentation. We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
@@
Prepending this string to participants’ prompt did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect. Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention. The GPT model still modified 59% of participant prompts. The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match). Conditional on any modification (any observations with TSR $<$ 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
A.2.2 Target Images
A.3 Subjects
Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once. We also prevented users from completing the task on mobile devices or tablets. Data was collected between December 12, 2023 and December 19, 2023. Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section C.5.1). The median time to complete our entire task, including a demographic survey, was 22 minutes. Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour. We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task. The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified. For subsequent checks, participants were required to retry until they demonstrated understanding.
A.4 Model Endpoints

### v6

@@ A.1 Model Setup
    Boundedness. Even for very large $x$, $Q<1$; quality cannot exceed perfection.
@@
    Diminishing returns. The marginal gain from effort, $Q_{x}=\theta se^{-\theta sx}$, declines as $x$ increases.
Prompting entails a linear cost function
$c(x)=kx,\qquad k>0,$
interpreted as the time or cognitive cost per unit of effort. The user’s utility from a given combination of model capacity, skill, and effort is
$U(\theta,s,x)=1-e^{-\theta sx}-kx.$
A.2 User Optimization
The user chooses effort $x\geq 0$ to maximize $U(\theta,s,x)$. The first-order condition is:
$\theta se^{-\theta sx}=k\;\;\Rightarrow\;\;x^{*}(\theta,s)=\frac{1}{\theta s}\ln\!\Bigl(\frac{\theta s}{k}\Bigr).$
An interior optimum requires $\theta s>k$, ensuring that $x^{(}\theta,s)>0$ and the logarithm is well defined. Intuitively, the user’s effective productivity $\theta s$ must exceed the marginal cost of effort $k$. This inequality is readily satisfied when prompting costs are small (i.e., $k\approx 0$).
Substituting the optimal effort into the quality function yields the user’s maximum attainable quality:
$Q^{*}(\theta,s)=1-\frac{k}{\theta s}.$ (6)
This closed-form expression will serve as the basis for all subsequent comparative statics and distributional analysis.
A.3 Comparative Statics
The partial derivatives of $Q^{*}$ are straightforward:
$\frac{\partial Q^{*}}{\partial\theta}=\frac{k}{\theta^{2}s}>0,\qquad\frac{\partial Q^{*}}{\partial s}=\frac{k}{\theta s^{2}}>0,\qquad\frac{\partial Q^{*}}{\partial k}=-\frac{1}{\theta s}<0.$
Hence, output quality increases in both model capacity and user skill, and decreases with the cost of effort. Improvements in model quality or prompting skill therefore unambiguously raise output quality, whereas higher marginal costs of effort reduce it.
A.4 Distributional Impact of Model Improvements (Overall Equalizing Effect)
We now examine how increases in model capacity $\theta$ reshape the distribution of optimal quality $Q^{*}$ across users. Because $Q^{*}$ is strictly increasing in skill $s$, dispersion in $Q^{*}$ reflects variation in skill, since $Q^{*}$ is a monotone transformation of $s$ (and hence of $G_{s}$). Let $G_{Q^{*}}(\pi)$ denote the inverse CDF of $Q^{*}$—that is, the quality achieved by the user at the $\pi$-th percentile of the quality distribution—and let $G_{s}(\pi)$ denote the inverse CDF of the skill distribution. Then:
$G_{Q^{*}}(\pi)=1-\frac{k}{\theta G_{s}(\pi)}.$
Proposition 1 (Equalizing Effect of Model Improvements).
For any strictly increasing skill distribution $G_{s}(\pi)$, we have
$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}<0.$
Proof.
Differentiating once gives
$\frac{\partial G_{Q^{*}}(\pi)}{\partial\theta}=\frac{k}{\theta^{2}G_{s}(\pi)}>0,$
so all users benefit from higher $\theta$. Differentiating again with respect to $\pi$,
$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{k}{\theta^{2}G_{s}^{2}(\pi)}\,G_{s}^{\prime}(\pi)<0,$
because $G_{s}^{\prime}(\pi)>0$. ∎
Model improvements therefore compress the distribution of output quality, reducing dispersion across users. Intuitively, greater model capacity raises quality throughout the distribution but shifts the lower portion upward more strongly in relative terms, since outputs at the top are already close to the performance ceiling.
A.5 Decomposing Total Improvements in Quality
The total change in optimal quality with respect to model capacity, $\theta$, can be expressed as a combination of a direct effect of the model itself and an indirect effect that operates through re-optimization of effort. To distinguish between these two direct and indirect channels, we introduce a counterfactual quality function that treats the model capacity determining effort and the model capacity entering the quality function as distinct parameters.
$Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))=1-\exp[-\theta_{2}s\,x^{*}(\theta_{1},s)]=1-\Bigl(\frac{k}{\theta_{1}s}\Bigr)^{\!\theta_{2}/\theta_{1}},$
which represents the quality achieved when a user employs the effort that would be optimal for model $\theta_{1}$ while using a model with capacity $\theta_{2}$. The model and prompting effects are obtained by differentiating this counterfactual quality function with respect to $\theta_{2}$ and $\theta_{1}$, respectively, and evaluating the results at $\theta_{1}=\theta_{2}=\theta$:
$M(\theta,s)$ $=\frac{\partial Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))}{\partial\theta_{2}}=\frac{k}{\theta^{2}s}\ln\!\Bigl(\frac{\theta s}{k}\Bigr)$
$P(\theta,s)$ $=\frac{\partial Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))}{\partial\theta_{1}}=\frac{k}{\theta^{2}s}\!\left[1-\ln\!\Bigl(\frac{\theta s}{k}\Bigr)\right].$
$M(\theta,s)$ captures the direct mechanical improvement in quality when model capacity increases but effort is held fixed. In contrast, $P(\theta,s)$, captures the behavioral response—the gain that results from users adjusting their effort once the model improves. By construction,
$\frac{dQ^{*}(\theta,s)}{d\theta}=M(\theta,s)+P(\theta,s).$
A.6 Distributional Implications of the Model and Prompting Effects
Although the decomposition in Appendix A.5 holds pointwise for each user, the distributional consequences of model and prompting effects require characterizing how these effects reshape the distribution of outcomes. Let $G_{Q^{cf}}(\pi)$ denote the inverse CDF of the counterfactual quality:
$G_{Q^{cf}}(\pi)=1-\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}.$
Proposition 2 (Model Effect: Equalizing Channel).
Holding prompting effort fixed, an increase in model capacity compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more. More formally,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}<0.$
Proof.
Differentiating with respect to $\theta_{2}$,
$\frac{\partial G_{Q^{cf}}(\pi)}{\partial\theta_{2}}=-\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}\frac{1}{\theta_{1}}\ln\!\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr).$
Differentiating again and setting $\theta_{1}=\theta_{2}=\theta$,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}=\frac{k}{\theta^{2}G_{s}^{2}(\pi)}G_{s}^{\prime}(\pi)\Bigl[\ln\!\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)+1\Bigr].$
For sufficiently small $k$, the bracketed term is negative, implying that the entire expression is negative. ∎
Proposition 3 (Prompting Effect: Expanding Channel).
Allowing users to re-optimize their prompting effort expands the distribution of output quality: performance differences across users widen as the upper portion of the distribution shifts further upward. More formally,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}>0.$
Proof.
Differentiating with respect to $\theta_{1}$,
$\frac{\partial G_{Q^{cf}}(\pi)}{\partial\theta_{1}}=\frac{\theta_{2}}{\theta_{1}^{2}}\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)^{\!\frac{\theta_{2}}{\theta_{1}}}\Bigl[\ln\!\Bigl(\frac{k}{\theta_{1}G_{s}(\pi)}\Bigr)+1\Bigr].$
Differentiating again and setting $\theta_{1}=\theta_{2}=\theta$,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}=-\frac{1}{\theta}\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)\frac{G_{s}^{\prime}(\pi)}{G_{s}(\pi)}\Bigl[\ln\!\Bigl(\frac{k}{\theta G_{s}(\pi)}\Bigr)+2\Bigr].$
For sufficiently small $k$, the term in brackets is negative, so the derivative is positive. ∎
A.7 Summary of Distributional Impacts
In this exponential-bounded specification, the distributional consequences of all effects are analytically unambiguous:
Effect Cross-partial sign Distributional interpretation
Model effect <0<0 Equalizing (Compression)
Prompting effect >0>0 Expanding (Dispersion)
Total effect <0<0 Equalizing (Compression)
The overall equalizing pattern established in Section A.4 thus arises because the compressive model channel dominates the dispersive prompting channel in magnitude. Intuitively, as model capacity increases, gains in output quality are largest at the lower end of the distribution, where outcomes remain far from the quality ceiling. Improvements at the upper end are more limited, leading to an overall compression in the distribution of performance. Behavioral adaptation through prompting has the opposite tendency, stretching the upper tail, but not enough to offset the dominant equalizing shift.
A.8 Unbounded Quality
We now consider a concave but unbounded quality specification,
$Q(\theta,s,x)=\log\!\big(1+(\theta+s)x\big)$
with the same linear cost $c(x)=kx$. This formulation preserves monotonicity and concavity in $x$, $\theta$, and $s$, but relaxes the upper bound on attainable quality. The user’s utility function is therefore
$U(\theta,s,x)=\log\!\big(1+(\theta+s)x\big)-kx$
User optimization.
The first-order condition is
$\frac{\theta+s}{1+(\theta+s)x}=k$
Assuming an interior solution ($\theta+s>k$), the optimal effort, quality, and utility are
$x^{*}(\theta,s)$ $=\frac{1}{k}-\frac{1}{\theta+s}$
$Q^{*}(\theta,s)$ $=\log\!\left(\frac{\theta+s}{k}\right)$
$U^{*}(\theta,s)$ $=\log\!\left(\frac{\theta+s}{k}\right)-1+\frac{k}{\theta+s}$
Comparative statics.
Optimal quality increases with both model capacity and user skill, and decreases with prompting cost:
$\frac{\partial Q^{*}}{\partial\theta}$ $=\frac{1}{\theta+s}>0\qquad\frac{\partial Q^{*}}{\partial s}=\frac{1}{\theta+s}>0\qquad\frac{\partial Q^{*}}{\partial k}=-\frac{1}{k}<0$
Distributional Impact of Model Improvements (Overall Equalizing Effect)
We now analyze how increases in model capacity $\theta$ reshape the distribution of optimal quality $Q^{*}$ across users. Let $G_{Q^{*}}(\pi)$ denote the inverse CDF of $Q^{*}$ and $G_{s}(\pi)$ the inverse CDF of skill. Then:
$G_{Q^{*}}(\pi)=\log\!\left(\frac{\theta+G_{s}(\pi)}{k}\right)$
Proposition 4 (Equalizing Effect of Model Improvements, Unbounded Case).
For any strictly increasing skill distribution $G_{s}(\pi)$,
$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{G_{s}^{\prime}(\pi)}{(\theta+G_{s}(\pi))^{2}}<0$
Proof.
Differentiating once gives
$\frac{\partial G_{Q^{*}}(\pi)}{\partial\theta}=\frac{1}{\theta+G_{s}(\pi)}>0$
so all users gain from higher $\theta$. Differentiating again with respect to $\pi$,
$\frac{\partial^{2}G_{Q^{*}}(\pi)}{\partial\theta\,\partial\pi}=-\frac{G_{s}^{\prime}(\pi)}{(\theta+G_{s}(\pi))^{2}}<0$
Model improvements therefore compress the quality distribution, raising performance throughout the distribution but proportionally more at lower quantiles—mirroring the equalizing effect found in the bounded model.
Counterfactual quality and decomposition.
As in the bounded model, we decompose total quality improvements into direct (model) and indirect (prompting) components using a counterfactual formulation:
$Q^{cf}(\theta_{2},s,x^{*}(\theta_{1},s))=\log\!\Big(1+(\theta_{2}+s)\,x^{*}(\theta_{1},s)\Big)$
where $\theta_{2}$ governs the model capacity in the quality function and $\theta_{1}$ determines the user’s optimal effort. Differentiating with respect to $\theta_{2}$ and $\theta_{1}$ and evaluating at $\theta_{1}=\theta_{2}=\theta$ gives
$M(\theta,s)=\frac{(\theta+s)-k}{(\theta+s)^{2}},\qquad P(\theta,s)=\frac{k}{(\theta+s)^{2}}$
so that the total effect of model improvement equals the sum of model and prompting effects: $\frac{dQ^{*}(\theta,s)}{d\theta}=M(\theta,s)+P(\theta,s)$ as before.
Distributional implications.
Let $G_{Q^{cf}}(\pi)$ denote the inverse CDF of counterfactual quality. It can be shown that $Q^{cf}$ is strictly increasing in $s$ if $k<\theta_{1}+s$, a condition needed for an interior solution. Thus:
$G_{Q^{cf}}(\pi;\theta_{1},\theta_{2})=\log\!\Bigg(1+\frac{\theta_{2}+G_{s}(\pi)}{k}-\frac{\theta_{2}+G_{s}(\pi)}{\theta_{1}+G_{s}(\pi)}\Bigg)$
Proposition 5 (Model Effect: Equalizing Channel, Unbounded Case).
Holding prompting effort fixed, an increase in model capacity compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more. Formally,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{2}\,\partial\pi}=\frac{2k-\big(\theta+G_{s}(\pi)\big)}{(\theta+G_{s}(\pi))^{3}}G_{s}^{\prime}(\pi)<0$
which is negative for sufficiently small $k$.
Proposition 6 (Prompting Effect: Equalizing Channel, Unbounded Case).
Allowing users to re-optimize their prompting effort compresses the distribution of output quality: the spread of outcomes narrows as performance at the lower end of the distribution rises proportionally more. Formally,
$\frac{\partial^{2}G_{Q^{cf}}(\pi)}{\partial\theta_{1}\,\partial\pi}=-\frac{2k}{(\theta+G_{s}(\pi))^{3}}G_{s}^{\prime}(\pi)<0.$
Appendix B Experiment Design Details
In this appendix, we provide details on the design of both experiments reported in the main text.
B.1 Randomization
For both experiments, we randomized participants across two dimensions: the target task (target image for the image replication experiment and target brief for the logo generation experiment) and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target item-model cells per experiment. We conducted a balance check after the conclusion of each experiment with a $\chi^{2}$ test across all cells. For the image replication experiment, we obtained $\chi^{2}=7.056,\;df=44,\;p>0.999$, thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups. For the logo generation experiment, we obtained $\chi^{2}=5.383,\;df=44,\;p>0.999$, thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups. In both experiments, participants were unaware of the underlying randomization.
B.1.1 Balance Checks
Table A.1 reports covariate balance checks for the image replication experiment across prior GPT usage, gender, and age. We find no statistically significant evidence of imbalance across any of these dimensions. Table A.2 reports the corresponding balance checks for the logo generation experiment. While one pairwise comparison yields a $p$-value below 0.05, this is not unexpected given that we conduct 18 balance tests across the two experiments.
Importantly, the statistically significant comparison arises in a contrast involving the DALL-E 3 (Revised) arm. As described in the main text, this arm is included primarily as a secondary, diagnostic condition to assess the consequences of automated system-side prompt rewriting, and it is not central to the paper’s primary treatment contrasts or replay-based identification strategy. We do not rely on direct comparisons between DALL-E 3 (Revised) and DALL-E 3 (Verbatim) for our headline results, and inspection of the balance table indicates that the deviation is driven by the DALL-E 3 (Revised) group rather than broad differences across the core arms of interest.
Finally, to assess whether this pattern could have been induced by sample restrictions or data processing, in Table A.3 we re-conduct the logo-generation balance checks on the full sample prior to applying any exclusions. The same low $p$-value persists, indicating that it is not created by filtering and is consistent with random variation in assignment rather than some systematic bias in our treatment of the data.
Table A.1: Covariate Balance Across All Three Arms (Image Replication Experiment)
Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
GPT Usage p = 0.745, $\chi^{2}$ = 1.948 (4) p = 0.654, $\chi^{2}$ = 2.449 (4) p = 0.629, $\chi^{2}$ = 2.587 (4)
None 215 0.329 186 0.305 186 0.296
Little 190 0.291 178 0.292 195 0.311
Some 179 0.274 173 0.284 170 0.271
A lot 70 0.110 72 0.118 76 0.121
NA/Missing 1 0.002 0 0.000 2 0.003
Gender p = 0.727, $\chi^{2}$ = 1.309 (3) p = 0.319, $\chi^{2}$ = 3.517 (3) p = 0.231, $\chi^{2}$ = 4.297 (3)
Female 354 0.541 332 0.545 341 0.544
Male 279 0.427 261 0.428 271 0.432
Other or prefer not to say 21 0.032 16 0.026 13 0.021
NA/Missing 1 0.002 0 0.000 4 0.006
Age Group p = 0.847, $\chi^{2}$ = 2.013 (5) p = 0.649, $\chi^{2}$ = 3.334 (5) p = 0.250, $\chi^{2}$ = 6.631 (5)
[18,26) 112 0.171 109 0.179 99 0.158
[26,32) 136 0.208 122 0.200 122 0.195
[32,39) 139 0.213 132 0.217 147 0.235
[39,49) 127 0.194 127 0.208 119 0.190
[49,83] 140 0.214 119 0.195 138 0.220
NA/Missing 1 0.002 0 0.000 4 0.006
  • Sample sizes: DALL-E 2 (D2) = 655, DALL-E 3 Revised (D3r) = 609, DALL-E 3 Verbatim (D3v) = 629.
    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm. Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
Table A.2: Covariate Balance Across All Three Arms (Logo Generation Experiment)
Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
GPT Usage p = 0.098, $\chi^{2}$ = 7.843 (4) p = 0.165, $\chi^{2}$ = 6.497 (4) p = 0.032, $\chi^{2}$ = 10.542 (4)
None 67 0.107 42 0.069 70 0.113
Little 117 0.186 123 0.201 100 0.162
Some 205 0.327 227 0.371 234 0.379
A lot 237 0.378 219 0.358 213 0.345
NA/Missing 2 0.003 1 0.002 0 0.000
Gender p = 0.901, $\chi^{2}$ = 0.581 (3) p = 0.563, $\chi^{2}$ = 2.046 (3) p = 0.417, $\chi^{2}$ = 2.838 (3)
Female 323 0.516 326 0.533 319 0.520
Male 297 0.474 277 0.453 293 0.480
Other / Prefer not to say 6 0.010 7 0.011 5 0.008
NA/Missing 2 0.003 2 0.003 0 0.000
Age Group p = 0.465, $\chi^{2}$ = 4.610 (5) p = 0.501, $\chi^{2}$ = 4.348 (5) p = 0.392, $\chi^{2}$ = 5.203 (5)
[18,29) 126 0.201 120 0.196 113 0.185
[29,35) 103 0.164 126 0.206 118 0.193
[35,43) 145 0.231 129 0.211 134 0.219
[43,52) 123 0.196 105 0.172 129 0.211
[52,87] 129 0.206 130 0.213 123 0.201
NA/Missing 2 0.003 2 0.003 0 0.000
  • Sample sizes: DALL-E 2 (D2) = 628, DALL-E 3 Revised (D3r) = 612, DALL-E 3 Verbatim (D3v) = 617.
    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm. Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
Table A.3: Covariate Balance Across All Three Arms (Logo Generation Experiment, Pre-filtering; N = 1924)
Covariate / Level N(D2) Prop(D2) N(D3r) Prop(D3r) N(D3v) Prop(D3v) D2 vs. D3r D2 vs. D3v D3r vs. D3v
p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df) p-val, $\chi^{2}$(df)
GPT Usage p = 0.134, $\chi^{2}$ = 7.044 (4) p = 0.267, $\chi^{2}$ = 5.203 (4) p = 0.034, $\chi^{2}$ = 10.441 (4)
None 67 0.103 44 0.070 73 0.114
Little 121 0.186 123 0.196 101 0.158
Some 212 0.326 235 0.375 240 0.374
A lot 248 0.382 223 0.356 227 0.354
NA/Missing 3 0.005 2 0.003 5 0.008
Gender p = 0.876, $\chi^{2}$ = 0.690 (3) p = 0.825, $\chi^{2}$ = 0.900 (3) p = 0.762, $\chi^{2}$ = 1.163 (3)
Female 330 0.508 332 0.530 332 0.518
Male 311 0.478 285 0.455 304 0.474
Other / Prefer not to say 7 0.011 7 0.011 5 0.008
NA/Missing 3 0.005 3 0.005 5 0.008
Age Group p = 0.568, $\chi^{2}$ = 3.873 (5) p = 0.877, $\chi^{2}$ = 1.793 (5) p = 0.636, $\chi^{2}$ = 3.418 (5)
[18,29) 130 0.200 122 0.195 118 0.184
[29,35) 111 0.171 130 0.208 121 0.189
[35,43) 147 0.226 130 0.208 140 0.218
[43,52) 128 0.197 109 0.174 133 0.207
[52,87] 132 0.203 133 0.212 129 0.201
NA/Missing 3 0.005 3 0.005 5 0.008
  • Sample sizes: DALL-E 2 (D2) = 651, DALL-E 3 Revised (D3r) = 627, DALL-E 3 Verbatim (D3v) = 646. Includes 7 users missing survey crosswalk.
    $N(\cdot)$ columns report the count of observations in each treatment arm belonging to that category, $\text{Prop}(\cdot)$ gives the proportion of respondents in that category for the arm. Chi-squared test columns show p-values and $\chi^{2}$ test statistics with degrees of freedom for each pairwise comparison.
B.2 Generative Models
For both experiments, we randomly assigned participants to 1 of 3 generative models:
  1. 1.
    DALL-E 2
  2. 2.
    DALL-E 3 (Verbatim) (sometimes simply referred to as DALL-E 3)
@@ B.2 Generative Models
    DALL-E 3 (Revised)
Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API. OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E 3 image-generating model. During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system. To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend text to the participant’s prompt instructing the GPT-4 model to refrain from modifying it before passing it forward to DALL-E 3. This prepended text was never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^11] See here and here for online documentation. We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The exact text we prepended to participant prompts is copied below:
@@
Prepending this text to participants’ prompts did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect. Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention. The GPT model still modified 59% of participant prompts. The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match). Conditional on any modification (any observations with TSR << 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
B.3 Model Endpoints